# Exponents of $2$-multiarrangements of three lines over fields of positive characteristic

### v2

@@ Abstract
Freeness of multiarrangements is extremely important in the study of freeness of hyperplane arrangements. In particular, the exponents of 2-multiarrangement play an important role in freeness of 3-arrangement. However, it is not easy to compute exponents in general.
Wakamiko (2007) determined bases and the exponents for multiarrangements of three lines over a field of characteristic zero. In this paper, we study the exponents of multiarrangements of three lines for the case of positive characteristic. We provide an effective algorithm for computing the exponents. Furthermore, we prove that the multiplicity lattice has plenty of symmetries. We also discuss the validity of bases constructed from binomial expansions in the case of positive characteristic.
@@ 1 Introduction
Freeness is one of central topics of study of arrangements. In particular, it is still an open problem whether the freeness of a hyperplane arrangement over a fixed field is determined solely by the structure of its intersection lattice (Terao conjecture). Freeness of multiarrangements plays an important role for understanding freenss of arrangements.
@@
Using this matrix, we have a useful criterion for freeness proved by K. Saito for simple arrangements [5, Theorem 1.8] and by Ziegler for multiarrangements [7, Theorem and Definition 8].
@@
In general, it is not easy to determine whether a multiarrangement is free or not. Every $2$-multiarrangement is known to be free [7, Corollary 7]. However, it is also not easy to compute a basis and exponents.
@@
where $\{x,y\}$ denote a basis for $V^{\ast}$. Let $\mu_{i}\coloneqq\mu(H_{i})$ for $i\in\{1,2,3\}$. Note that $\mu_{1},\mu_{2}$ and, $\mu_{3}$ have the same role by symmetry of $H_{1}$, $H_{2}$, and $H_{3}$.
@@
Theorem 1.2 (Wakamiko [6, Theorem 1.5]).
@@
    where $\lfloor x\rfloor=\max{\Set{n\in\mathbb{Z}}{n\leq x}}$, and $\lceil x\rceil=\min{\Set{n\in\mathbb{Z}}{x\leq n}}$. In this case, a basis for $D(\mathcal{A},\mu)$ can be expressed by using generalized binomial coefficients (See [6] for the description).
@@
Feigin, Wang, and Yoshinaga [3, Theorem 2.3] constructed bases for $D(\mathcal{A},\mu)$ with integral expressions for balanced cases.
Let $\mathit{\Delta}(\mu)$ denote the absolute value of the difference between exponents of $(\mathcal{A},\mu)$. By Theorem 1.2, the exponents can be divided into unbalanced and balanced cases as follows.
@@
form a basis for $D(\mathcal{A},\mu)$ (See [6, Example 4.1]) since
@@
As shown in Table 1, the multiplicity lattice has plenty of symmetries. In Section 4, we prove of the periodicity (Theorem 4.2) and symmetry (Theorem 4.5) by constructing correspondences between bases.
@@ 3 Exponents
In this section, we suppose that $\mathbb{F}=\mathbb{F}_{p}$ with $p$ prime and $\mathcal{A}$ is a $2$-arrangement over $\mathbb{F}_{p}$ consisting of three lines. When $\mu$ is unbalanced, by Theorem 2.4,
@@ 3.1 Centers of connected components
Conversely, suppose that $\mu$ belongs to the right hand side. Then $\mathit{\Delta}(\mu)\geq 1$ since $\mu\in\mathit{\Lambda}^{\textup{bal}}_{\textup{odd}}$. Therefore $\mu$ belongs to some connected component of $\mathit{\Lambda}^{\textup{bal}}_{>0}$. Namely, by Theorem 3.1, there exist $m\in\mathbb{Z}_{\geq 0}$ and $\zeta\in p^{m}\mathit{\Lambda}_{\textup{odd}}^{\textup{bal}}$ such that $\mu\in B(\zeta,p^{m})$. This implies $\mu\in B(p^{m}\mathit{\Lambda}^{\textup{bal}}_{\textup{odd}},\,p^{m})$. From the assumption, we have $m=0$ and hence $\mu=\zeta\in Z(1)$. ∎
@@ 4 Periodicity and symmetry of exponents
In this section, we suppose that $\mathbb{F}=\mathbb{F}_{p}$ with $p$ prime and $(\mathcal{A},\mu)$ is defined by
$Q(\mathcal{A},\mu)=x^{\mu_{1}}y^{\mu_{2}}(x+y)^{\mu_{3}}.$ (4.1)
@@ 4.2 Symmetry
$\mathit{\Lambda}_{\leq p^{d}}\coloneqq\Set{(\mu_{1},\mu_{2},\mu_{3})\in\mathit{\Lambda}}{\mu_{1},\mu_{2},\mu_{3}\leq p^{d}}.$ (4.22)
@@ 5 The basis with binomial coefficients
In some cases, a basis for $D(\mathcal{A},\mu)$ can be easily constructed using binomial coefficients, as in the unbalanced case for $\operatorname{char}{\mathbb{F}}=0$.
Suppose that $\operatorname{char}{\mathbb{F}}=p>0$. For $m\in\mathbb{Z}_{>0}$, we denote the base $p$ expansions of $m$ as
@@
We will now consider to explicitly represent the set $\mathit{\Gamma}(m)$. A non-empty subset $\mathcal{L}$ of $\mathit{\Lambda}(m)$ is a lower set of $\mathit{\Lambda}(m)$ if $\mu\subseteq\nu$ for $\mu\in\mathit{\Lambda}(m)$ and $\nu\in\mathcal{L}$ implies $\mu\in\mathcal{L}$. We denote by $\mathcal{L}(S)$ the lower set of $\mathit{\Lambda}(m)$ generated by $S\subseteq\mathit{\Lambda}(m)$.
@@
A non-empty subset $\mathcal{U}$ of $\mathit{\Lambda}(m)$ is an upper set of $\mathit{\Lambda}(m)$ if $\mu\supseteq\nu$ for $\mu\in\mathit{\Lambda}(m)$ and $\nu\in\mathcal{U}$ implies $\mu\in\mathcal{U}$. We denote by $\mathcal{U}(B)$ the upper set of $\mathit{\Lambda}(m)$ generated by $B\subseteq\mathit{\Lambda}(m)$.
@@
Table 3 shows the lowest degree of $D(\mathcal{A},\mu)$ for each $\mu\in\mathit{\Lambda}(m)$. A multiplicity $\mu$ corresponding to the cell colored in light gray in Table 3 belongs to $\mathit{\Gamma}(m)$. Moreover, ()($\cdot$)( ⋅ ) denotes that the multiplicity is the center of the connected component.
@@
To prove the statement, since $\mathit{\Delta}|_{C_{\kappa,t}}$ is unimordal, we only show that $\mathit{\Delta}(\mu)<\mathit{\Delta}(\kappa)$ for any $\nu\in C_{\kappa,t}$ adjacent to $\kappa\in S(m)$. If $\mu\subseteq\kappa$, then $\nu\in\mathit{\Gamma}(m)$. Since $g+g^{\prime}\geq m+1$ by (5.38), it satisfies $\mu_{1}+\mu_{2}\geq m$. Hence Theorem 5.6 implies that

### v3

@@ Abstract
Wakamiko determined bases and the exponents for multiarrangements of three lines over a field of characteristic zero. In this paper, we study the exponents of multiarrangements of three lines for the case of positive characteristic. We provide an effective algorithm for computing the exponents. Furthermore, we prove that the multiplicity lattice has plenty of symmetries. We also discuss the validity of bases constructed from binomial expansions in the case of positive characteristic.
@@ 1 Introduction
Freeness is one of central topics of study of arrangements. In particular, it is still an open problem whether the freeness of a hyperplane arrangement over a fixed field is determined solely by the structure of its intersection lattice (Terao conjecture [7, Problem 1]). Freeness of multiarrangements plays an important role for understanding freenss of arrangements [9, 10, 3].
@@
Using this matrix, we have a useful criterion for freeness proved by K. Saito for simple arrangements [6, Theorem 1.8] and by Ziegler for multiarrangements [11, Theorem and Definition 8].
@@
In general, it is not easy to determine whether a multiarrangement is free or not. Every $2$-multiarrangement is known to be free [11, Corollary 7]. However, it is also not easy to compute a basis and exponents.
@@
where $\{x,y\}$ denotes a basis for $V^{\ast}$. Let $\mu_{i}\coloneqq\mu(H_{i})$ for $i\in\{1,2,3\}$. Note that $\mu_{1},\mu_{2}$ and, $\mu_{3}$ have the same role by symmetry of $H_{1}$, $H_{2}$, and $H_{3}$.
@@
Theorem 1.2 (Wakamiko [8, Theorem 1.5]).
@@
    where $\lfloor x\rfloor=\max{\Set{n\in\mathbb{Z}}{n\leq x}}$, and $\lceil x\rceil=\min{\Set{n\in\mathbb{Z}}{x\leq n}}$. In this case, a basis for $D(\mathcal{A},\mu)$ can be expressed using generalized binomial coefficients (see [8] for the description).
@@
Feigin, Wang, and Yoshinaga [4, Theorem 2.3] constructed bases for $D(\mathcal{A},\mu)$ with integral expressions for balanced cases.
Let $\mathit{\Delta}(\mu)$ denote the absolute value of the difference between the exponents of $(\mathcal{A},\mu)$. By Theorem 1.2, the exponents can be divided into unbalanced and balanced cases as follows.
@@
form a basis for $D(\mathcal{A},\mu)$ (See [8, Example 4.1]) since
@@
As shown in Table 1, the multiplicity lattice has a lot of symmetries. In Section 4, we prove the periodicity (Theorem 4.2) and symmetry (Theorem 4.5) by constructing correspondences between bases.
@@ 3 Exponents
In the rest of this paper, let $(\mathcal{A},\mu)$ be the $2$-multiarrangement over a finite field $\mathbb{F}_{p}$ with $p$ prime defined by
$Q(\mathcal{A},\mu)=x^{\mu_{1}}y^{\mu_{2}}(x+y)^{\mu_{3}}.$ (3.1)
When $\mu$ is unbalanced, by Theorem 2.4,
@@ 3.1 Centers of connected components
Conversely, suppose that $\mu$ belongs to the right-hand side. Then $\mathit{\Delta}(\mu)\geq 1$ since $\mu\in\mathit{\Lambda}^{\textup{bal}}_{\textup{odd}}$. Therefore $\mu$ belongs to some connected component of $\mathit{\Lambda}^{\textup{bal}}_{>0}$. Namely, by Theorem 3.1, there exist $m\in\mathbb{Z}_{\geq 0}$ and $\zeta\in p^{m}\mathit{\Lambda}_{\textup{odd}}^{\textup{bal}}$ such that $\mu\in B(\zeta,p^{m})$. This implies $\mu\in B(p^{m}\mathit{\Lambda}^{\textup{bal}}_{\textup{odd}},\,p^{m})$. From the assumption, we have $m=0$ and hence $\mu=\zeta\in Z(1)$. ∎
@@ 4.2 Symmetry
$\mathit{\Lambda}_{\leq p^{d}}\coloneqq\Set{(\mu_{1},\mu_{2},\mu_{3})\in\mathit{\Lambda}}{\mu_{1},\mu_{2},\mu_{3}\leq p^{d}}$ (4.21)
@@ 5 The basis with binomial coefficients
For $m\in\mathbb{Z}_{>0}$, we express the base-$p$ expansion of $m$ as
@@
We will now consider to explicitly represent the set $\mathit{\Gamma}(m)$. A non-empty subset $\mathcal{L}$ of $\mathit{\Lambda}(m)$ is a lower set of $\mathit{\Lambda}(m)$ if $\mu\subseteq\nu$ for $\mu\in\mathit{\Lambda}(m)$ and $\nu\in\mathcal{L}$ implies $\mu\in\mathcal{L}$. Let $\mathcal{L}(S)$ denote the lower set of $\mathit{\Lambda}(m)$ generated by $S\subseteq\mathit{\Lambda}(m)$.
@@
A non-empty subset $\mathcal{U}$ of $\mathit{\Lambda}(m)$ is an upper set of $\mathit{\Lambda}(m)$ if $\mu\supseteq\nu$ for $\mu\in\mathit{\Lambda}(m)$ and $\nu\in\mathcal{U}$ implies $\mu\in\mathcal{U}$. Let $\mathcal{U}(B)$ denote the upper set of $\mathit{\Lambda}(m)$ generated by $B\subseteq\mathit{\Lambda}(m)$.
@@
Table 3 shows the lowest degree of $D(\mathcal{A},\mu)$ for each $\mu\in\mathit{\Lambda}(m)$, where the shaded area denotes $\mathit{\Gamma}(m)$ and the brackets indicate the centers of connected components.
@@
To prove the statement, since $\mathit{\Delta}|_{C_{\kappa,t}}$ is unimordal, we only show that $\mathit{\Delta}(\mu)<\mathit{\Delta}(\kappa)$ for any $\nu\in C_{\kappa,t}$ adjacent to $\kappa\in S(m)$.
If $\mu\subseteq\kappa$, then $\nu\in\mathit{\Gamma}(m)$. Since $g+g^{\prime}\geq m+1$ by (5.38), it satisfies $\mu_{1}+\mu_{2}\geq m$. Hence Theorem 5.6 implies that